The sum of 3 consecutive integers is 549. What is the middle number
Answer by chibisan(120)

1st number : x
middle number : x+1
3rd number : x+2
x+x+1+x+2 = 549
3x + 3 = 549
3x = 546
x = 186
middle number : 186+1 = 187.